A BLOCK WEIGHING 10KG IS AT REST ON HORIZONTAL TABLE THE COEFFICIENT FRICTION BETWEEN BLOCK AND TABLE IS 0.5 IF A FORCE ACTS DOWNWARDS AT 60 DEGREE WITH THE HORIZONTAL , HOW LARGE CAN IT TO BE WITHOUT CAUSING THE BLOCK TO MOVE

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
a = 0.5 * 10 = 5 m/sec^2
 
Now
 
F = ma = 10 * 5 = 50N
 
Hence it can be any force till you apply 50N without causing the block to move